OpenStudy (anonymous):

Find S12 for the series 1 + 1.5 + 2.25 + 3.375 +...

Parth (parthkohli):

A geometric series with common ratio 3/2.

OpenStudy (anonymous):

I got 86.49 as the 12 number in the series. Is that correct?

OpenStudy (triciaal):

verify that the question is not asking for the 12th term it is asking for the sum of the first 12 terms.

Parth (parthkohli):

Yes, but if you have to find the sum, you use\[S (n) = a\times\dfrac{ r^n-1 }{r -1}\]

OpenStudy (anonymous):

okay, I was just wondering if the number that I got for the 12th term was correct.

OpenStudy (triciaal):

so you wanted T12 ok T12 = ar^n-1 a= 1, n = 12, r = 3/2 T12 = (1)(3/2)^(12-1) = (3/2)^11 =

OpenStudy (anonymous):

I got 258.49 as my over all answer, Is this correct?

OpenStudy (anonymous):

IS 258.49 correct for the sum of the terms

OpenStudy (triciaal):